what is negative 3 over 5 multiplied by a number equal to 30
Found 2 solutions by Theo, checkley77:
Answer by Theo(13342)

-3/5 * x = 30

multiply both sides by 5 to get:

-3*x = 30*5

divide both sides by -3 to get:

x = 30*5/(-3)

solve for x to get:

x = -50

answer should be x = -50

plug into original equation to get:

-3/5 * x = 30 becomes:

-3/5 * -50 = 30

simplify to get:

-3 * -10 = 30

simplify further to get:

30 = 30

answer of -50 looks good.


Answer by checkley77(12844)

-3/5*X=30
-3X/5=30
-3X=5*30
-3X=150
X=150/-3
X=-50 ANS.
PROOF:
-3/5*-50=30
150/5=30
30=30